OpenStudy (anonymous):

Solve 7 and four-ninths minus 3 and two-thirds. Show your work

OpenStudy (anonymous):

if you mean 4 1/5 - 2 2/3, first you change them both into mixed fractions: 21/5 - 8/3 you change both denominators to 15: (3)21/5 - (5)8/3 = 63/15 - 40/15 and now you subtract: 63/15 - 40/15 = 23/15 last but not least you simplify: 23/15 = 1 8/15 final answer: 1 8/15

OpenStudy (anonymous):

two ninths

OpenStudy (anonymous):

7 4/9 - 3 2/3

OpenStudy (anonymous):

hold on im sowing you how to do the promblem

OpenStudy (anonymous):

problem

OpenStudy (anonymous):

k

OpenStudy (anonymous):

one ninth that's the answer

OpenStudy (anonymous):

hold on right quick

OpenStudy (anonymous):

www.webanswers.com › Categories › Education › Homework Questions go to the site

OpenStudy (jojo4eva):

1/9 is the answer